jhonyy9 (jhonyy9):

how is possible calculate every prim numbers from 2 till ... the end - how,what is this formula ... so if may be this exist today,anywhere in present ? - or this is very very difficile to create it ?

OpenStudy (amistre64):

prime number fascination has been around a long long time :) a cant recall any of the formulas they tried but its a fun research topic

OpenStudy (a_clan):

there are infinite number of prime numbers. no formula can calculate all of them. but computer programs can help

jhonyy9 (jhonyy9):

if we multiple for example : 1x2=2+1=3 2x2+1=5 --- 2x3+1=7 --- 2x5+1=11 ---

jhonyy9 (jhonyy9):

or 2x3x5+1=31

OpenStudy (a_clan):

2*7+1 =15 (not prime) I cannot say how reliable this method is. It does not seem to follow a fixed pattern

jhonyy9 (jhonyy9):

yes always exist exceptions

OpenStudy (amistre64):

kid in algebra class had the idea that if it ends in a 3 its prime; I was like really? so when they got to 33 I just smiled :)

OpenStudy (anonymous):

if you could do such a thing a) you would be very famous b) no one's atm cards would be worth the plastic they are made out of

OpenStudy (anonymous):

what is known is how the prime numbers are distributed, not how to find them. turns out that for an integer n the number of primes less than n is around\[ \frac{n}{\ln(n)}\] it has been known since Euclid (300bc) that there are an infinite number of primes

jhonyy9 (jhonyy9):

- if we multiple the primes what we know and to this sum we add one we will get prim ? - like 2x3x5+1=31 - or like 2x3x5x7+1=211 - this is true ?

OpenStudy (anonymous):

That is the basis for Euclid's proof ie from a supposedly complete list of primes, you could construct another one (they weren't really into infinity, the Greeks).

jhonyy9 (jhonyy9):

- the reality for what i have asked this question is because i have one formula with what i can calculate every primes from 5 till 101 - what i have probed - - how much is it interesting ?

OpenStudy (anonymous):

What would be interesting is discovery of any statements that can be made about primes in general (or even identifiable groups of them). eg For n>1, there is a prime between n and 2n. It is generally thought that no formula exists that will produce all the primes.

jhonyy9 (jhonyy9):

- ok so from your last words i need understanding that if my formula ,,work" how i have wrote till what i have probed it - is it very very usefuly - important - ?

OpenStudy (anonymous):

Unfortunately not, maybe you should take a look at http://en.wikipedia.org/wiki/Prime_formula

jhonyy9 (jhonyy9):

thank you